OpenStudy (anonymous):

over the set of real numbers, what is the domain of the composite function g(f(x))? f(x)=3-x,g(y)=square root of y g(f(x))=square root of 3-x

OpenStudy (anonymous):

hey i got \[x \le3 \]

OpenStudy (mathstudent55):

\( g(y) = \sqrt{y} \) \(f(x) = 3 - x\) \(g(f(x)) = g(3 - x)\) \(g(f(x)) = \sqrt{3 - x} \) The domain is the set of values that will make the radicand non-negative: \(3 - x \ge 0\) \(-x \ge -3\) \(x \le 3 \) You are correct.

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

wlcm